OpenStudy (anonymous):

Can someone please help me? I'm stuck. y=-1/4x+1/2-3 How do I cancel out the fractions?

OpenStudy (anonymous):

First, the purpose of cancelling in fractions is to simplify calculations involving them. For example, why multiply 68*34 when you can do 2*1. Second, the basic rule is one of division. Find the largest number you can divide both the top and bottom by (the GCF) and make sure you use the same value on both the top and bottom. For example, take 9/12. The GCF is 3, so divide top and bottom by 3: you get 3/4. Isn't that easier to work with? This rule can also be used in equations and algebraic expressions, etc. but this is the basic method. Keep in mind though, you can't subtract from top and bottom, 3/4 doesn't equal 1/2.

OpenStudy (anonymous):

Alright... But what do I do, how do I cancel it out? I can factor them anymore, now how do I cancel them out...? You didn't really help much----

OpenStudy (anonymous):

can't*

OpenStudy (anonymous):

Sorry..... Ugh, i hate doing problems. Try going to youtube, or google may help :) Sorry i suck at doing problems ill probably get it wrong :(

OpenStudy (anonymous):

.... Alright, thanks(?)

OpenStudy (anonymous):

;-; sorry..

OpenStudy (anonymous):

It's fine. @amistre64

OpenStudy (texaschic101):

to get rid of the fractions and turn them into whole numbers, you must multiply the entire equation by the LCD, which in this case is 4

OpenStudy (anonymous):

get rid of the fraction by multiplying them to their reciprocal

OpenStudy (texaschic101):

Thats if you want it in standard form...if you want it in slope intercept form, there is no reason to cancel the fractions

OpenStudy (anonymous):

I want standard form, so the whole entire equation. My answer should be-- y=-1x-1?

OpenStudy (texaschic101):

y = -1/4x + 1/2 - 3 y = -1/4x + 1/2 - 6/2 y = -1/4x - 5/2 (multiply by 4) 4y = -x - 10 x - 4y = -10 (standard form is Ax + By = C) I do not know how you got y = -x - 1

OpenStudy (anonymous):

Where'd you get 6/2 is my question---